OpenStudy (anonymous):

There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B. Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically

OpenStudy (boldjon):

Part A: \[f(x)=30*0.2x\]and\[f(x)=45+3x\] Part B: \[f(x)=30+0.2(5)\]\[f(x)=45+3(5)\]For Part B solve them. Part C: No idea. Ask someone else.